The scale on a map indicates that 1.5 centimeters represent 200 miles .if the distance on the map between Norfolk, Virginia , and Atlanta, Georgia , measures 2.4 centimeters , how many miles apart are the cities ?

Answer by mananth(16946)

direct proportion
1.5/2.4 = 200/x
x=2.4*200/1.5
x=320 miles
...
m.ananth@hotmail.ca